OpenStudy (baby456):

Gregg mutiplied a number by 5 subtract 6 from the product and got 29. what was his orginal number?

OpenStudy (anonymous):

29+6=35 35/5=7 that was simple.

OpenStudy (anonymous):

Check answer. 7*5=35-6=29